OpenStudy (anonymous):

I GIVE MEDALS Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (anonymous):

@MAEMAEHOCKEY

OpenStudy (anonymous):

The first step is to find the slope. The formula for slope is: \[m=(y _{2}-y _{1})/(x _{2}-x _{1})\] Where your two points correspond so: \[(x _{1},y _{1})=(-3,2)\] and \[(x _{2},y _{2})=(2,1)\] Can you find the slope now?

OpenStudy (mathmale):

The point-slope form of the equation of a straight line is\[y-y _{1}=m(x-x _{1})\] Follow @Alphabet_Sam 's advice and find the slope, m, of the line first.

OpenStudy (mathmale):

Start with what you know. I'm sure y ou must have seen the slope formula before. Exactly what does "slope of a straight line" mean to you?

OpenStudy (mathmale):

You are given two points and are told that a straight line runs thru both points. Your job is to find the equation of this line. Finding the slope of this line is the first step. Alphabet_Sam The first step is to find the slope. The formula for slope is: m=(y2−y1)/(x2−x1)

OpenStudy (mathmale):

that's the easy way out. Sorry, but OpenStudy is set up to help you learn and understand how to solve your own problems. No direct answers. It'd be faster if you'd please follow my suggestions and those of @Alphabet_Sam .

OpenStudy (dayakar):

equation of the line passing through (x1, y1) =(-3 ,2) and (x2 ,y2)= (2 ,1) is y -y1 = (y2 - y1)/(x2 - x1) *(x- x1) substite the values and get the equation